OpenStudy (anonymous):

Suppose that sec(B) = - 8⁄3 and that 180o < B < 270o. Find sin(B).

OpenStudy (anonymous):

cos B=-3/8 sin b=sqrt(1-(cos x)^2) sin b=sqrt(1-9/64) sin b=sqrt(55/64) since sin is negative in thord quadrant so sin b=-sqrt(55)/8.

OpenStudy (anonymous):

great!